OpenStudy (anonymous):

A package contains 5 candy canes, 3 of which are cracked. If 2 are selected, find the probability of getting no cracked candy canes

OpenStudy (anonymous):

2/5

OpenStudy (anonymous):

i should think not

OpenStudy (anonymous):

\[\frac{2}{5}\times \frac{1}{4}\]